OpenStudy (pphalke):

There were 375 tickets sold for a basketball game. The activity cardholders cost $1.25 and the non-cardholders tickets cost $2.25. The total amount of money collected was $1315.75. How many of each kind of ticket were sold?

OpenStudy (xguardians):

1.25x + 2.25y = 1315.75 x + y = 375

OpenStudy (pphalke):

how do i go about solving this?

OpenStudy (kimbrown415):

I think I know how to solve this one but even if you multiply 375 by $2.25 it doesnt work out the dollar amount is too low

OpenStudy (pphalke):

how @Kimbrown415

OpenStudy (xguardians):

No wonder my work wasn't making sense, if 375 tickets were sold at the premium price of $2.25, that doesn't even equal the total cost of $1315.75

OpenStudy (kimbrown415):

@xGuardians gave the formula . yes guardian it doesnt add up . Are you sure the problem is correct on screen?

OpenStudy (pphalke):

sorry my mistake its 735 tickets my mistake

OpenStudy (beth12345):

loool... dyslexia, that happens to me a lot too :p

OpenStudy (xguardians):

1.25x + 2.25y = 1315.75 Solve for a variable, in this case, y. y = 735 - x Substitute into other equation 1.25x + 2.25(735-x) = 1315.75 1.25x + 1653.75 - 2.25x = 1315.75 Solve for x 1.25x - 2.25x = -338 -1x = -338 x = 338 Therefore y = 735 - 338 = 397 Cardholders = 338 Nonecardholders = 397

OpenStudy (beth12345):

how did you get to the y=735-x part